OpenStudy (anonymous):

Erika ate two slices of a pizza with a 20" diameter. The pizza was sliced into 8 congruent slices. What was the area of the pizza that Erika ate?

OpenStudy (danjs):

This pizza has the same area as the last problem... 100*pi She ate 2 out of 8 slices, or 1/4 the pizza

OpenStudy (danjs):

1/4th of (100pi) is 25pi square inches that she ate

OpenStudy (anonymous):

@DanJS so what would the length of the crust of each pizza be? Is it 3.125?

OpenStudy (danjs):

The length of the crust is found by the circumference of the pizza circle. 2*pi*radius = 2*pi*10 = 20*pi The whole pizza is 20pi around, each piece is 1/8 th of that